# long-var-wrap = dv

### v1

@@ Abstract
Interactions between active particles may be non-reciprocal, breaking action-reaction symmetry and leading to novel physics not seen in equilibrium systems. In this work, we study the phase diagram for the non-reciprocal Cahn-Hilliard (NRCH) model with different types of symmetry—discrete and continuous—and examine the effects of non-reciprocity on phase separation. The models are analyzed by drawing the full linear stability diagrams, indicating the onset of phase separation, allowing us to demonstrate how non-reciprocity gives rise to out-of-equilibrium steady states with non-zero currents. These observations are robust to noise and valid in a large part of parameter space. By describing the effects of symmetry on conserved, non-reciprocal models, this work adds to our understanding of the rich phase-behavior of such models.
@@ II The Model
where summation is implied. The total free energy reads $F=\int\mathrm{d}\bm{x}\,\mathcal{F}$. The two models are defined by their potential $\mathcal{V}(\varphi)$. The potential of the $\mathrm{SO}(2)$ model
@@
If $r\geq 0$, the potential has one global minimum at $\varphi_{a}=0$, whereas for $r<0$ the potential takes the characteristic Mexican hat shape illustrated in Figure 1. The potential is a function of $\varphi_{a}\varphi_{a}$ only, which is a construction that is invariant under rotations and reflections (e.g. $\varphi_{1}\rightarrow-\varphi_{1}$) in the field space. These transformations form the Lie group $\mathrm{O}(2)$. The dynamics arising from this potential corresponds to Model B in the classification of Halperin and Hohenberg [53]. The symmetry is reduced to $\mathrm{SO}(2)$ with the introduction of a non-reciprocal term $\alpha\epsilon_{ab}\varphi_{b}$ in the equations of motion (see below).
@@
which is not invariant under continuous rotations, as illustrated in Figure 1. However, $\mathcal{V}_{\mathrm{d}}$ is invariant under operations in $\mathrm{D}_{4}$, the dihedral group of order 8 corresponding to the symmetries of a square. These operations are rotations with $n\pi/2$ radians in addition to reflections, e.g. $\varphi_{1}\rightarrow-\varphi_{1}$. This is a discrete group, similar to the symmetry group of the single-field Cahn-Hilliard $\mathrm{O}(1)\cong\mathrm{D}_{2}$, and in the same way gives rise to a phase separation with a well-defined domain wall. The introduction of a non-reciprocal term restricts the symmetry to $\mathrm{C}_{4}$, namely, the cyclic group of order 4.
Figure 2: The linear phase diagram of the model with continuous symmetry. The exceptional plane is highlighted in purple, and the planes between the stable, type-I unstable, and type-II unstable are highlighted in black, blue, and green, respectively. The red line is the intersection of all these surfaces.
@@ III Linear Stability
We now construct phase diagrams for the two cases that represent different symmetries by using linear stability analysis. We study the response of the system to small perturbations around a homogeneous state, represented by $\bar{\varphi}_{a}$. Inserting $\varphi_{a}(\bm{x},t)=\bar{\varphi}_{a}+\delta\varphi_{a}(\bm{x},t)$ into the equations of motion, we obtain the leading order contribution in Fourier space as follows
@@ III.1 Continuous Symmetry
The base of this diagram is the $\alpha=0$ surface. We choose $\alpha>0$, without loss of generality. At $\alpha=u\bar{\varphi}^{2}$, we have $\xi^{-2}_{+}=\xi^{-2}_{-}$. In reciprocal systems, this indicates a degeneracy where two independent eigenmodes have the same eigenvalue. When we deal with non-reciprocal systems, and thus non-Hermitian matrices, we are no longer guaranteed that the eigenvectors of $M_{ab}$ span the full vector space of perturbations. At this exceptional point (EP) [36, 56], the eigenvalues coincide because the two eigenmodes $\delta\varphi_{1}$ and $\delta\varphi_{2}$ align, rendering the matrix $M_{ab}$ defective. As there are three free parameters and the exceptional points are given by one equation, namely, $\alpha=u\bar{\varphi}^{2}$, the full set of exceptional points forms a 2D surface, which we denote as the exceptional plane. This is illustrated in purple in the Figure 2. Above the exceptional plane, the two eigenvalues become complex, and the perturbations will no longer decay or diverge monotonically: they exhibit oscillatory behavior. Here, we have $\mathrm{Re}\left(\xi_{\pm}^{-2}\right)=r+2u\bar{\varphi}^{2}$, and, therefore, both modes transition from stable to unstable at the same surface, defined via $r=-2u\bar{\varphi}^{2}$. The intersection between the critical and the exceptional planes can be called critical exceptional points (CEPs) [45], also corresponding to a Hopf bifurcation in the terminology of dynamical systems [57]. These CEPs are parameterized as follows
@@
Figure 3: The four non-trivial slices of the full three-dimensional phase diagram shown in Figure 2. Panel (a) corresponds to the equilibrium case $\alpha=0$, while panels (b)-(d) illustrate three orthogonal slices for finite $\alpha$, at constant $\alpha$, $\bar{\varphi}$, and $r$, respectively. Light blue and hashed areas mark the type-I and type-II unstable regions. Black, blue dash-dotted, and green lines separate stable and type-I, type-I and type-II, and stable and type-II, respectively. The purple dashed line marks the exceptional points, and the red dot marks the critical exceptional points where all the lines intersect.
To better present the phase diagram, we now examine various two-dimensional slices that can be obtained by fixing one of the parameters. Setting $\alpha=0$ gives the diagram of Model B for $n=2$ fields, with $\xi_{\pm}^{-2}=r+(2\pm 1)u\bar{\varphi}^{2}$. This is illustrated in Figure 3(a), and is the base of the diagram in Figure 2. At $r=\bar{\varphi}=0$, the correlation lengths diverge. This is a critical point corresponding to the onset of phase separation. Starting in the white region of the diagram, where $u\bar{\varphi}^{2}>|r|$, we have $\xi_{\pm}^{-2}>0$, and thus both modes are stable. Upon further lowering $\bar{\varphi}$, we reach a line in the diagram at $u\bar{\varphi}^{2}=|r|$ – shown as a black line in Figure 2 – where $\xi^{-2}_{-}=0$ and the corresponding eigenmode transitions from stable to unstable. Therefore, perturbations in this direction in $\varphi$-space ($\delta\varphi_{2}$ in this case) will grow exponentially.
Looking back at the Mexican hat potential illustrated in Figure 1, this corresponds to a point inside the ring of minima, and perturbations that would correspond to the Goldstone mode in a spontaneous symmetry breaking scenario (i.e. the orthogonal to the radial direction). The mode corresponding to perturbations in the radial direction ($\delta\varphi_{1}$ in this case) remains stable. This holds in the light blue area labeled “type-I unstable”, where there is one unstable mode. At $u\bar{\varphi}^{2}=|r|/3$, illustrated by a blue dash-dotted line, the second mode also transitions from stable to unstable. The area where both modes are unstable is highlighted as dark blue and hashed, with the label “type-II unstable”.
@@
Figure 4: The linear phase diagram of the model with discrete symmetry, scaled by $|r|$. Light blue and hashed areas mark the type-I and type-II unstable regions. Black, blue, and green planes separate stable and type-I, type-I and type-II, and stable and type-II regions, respectively. The exceptional plane is highlighted in purple, and the red line marks the critical exceptional points where all the planes intersect.
@@ III.2 Discrete Symmetry
Since the eigenvalues depend on both $\bar{\varphi}_{1}^{2}$ and $\bar{\varphi}_{2}^{2}$, the full phase diagram requires another dimension to visualize. For this model, we focus on a 3D slice of the full 4D diagram as shown in Figure 4. This cross section is related to any other 3D slice via a choice of scaling, which in this case is chosen to be $|r|$. The inverse-squared correlation lengths defined via $\xi_{\pm}^{-2}=\lambda_{\pm}(q\rightarrow 0)$ as above are then given as follows
@@
Figure 6: Constant $\alpha$ slices of the phase diagram illustrated in Figure 4. The gray dashed lines represent the line connecting the minima of the potential illustrated in Figure 1(a), and form a square centered at the origin and with side lengths $\sqrt{u/|r|}\bar{\varphi}=\sqrt{2}$. Light blue and hashed areas mark the type-I and type-II unstable regions. Black, blue dash-dotted, and green lines separate stable and type-I, type-I and type-II, and stable and type-II, respectively. The purple line marks the exceptional points, and the red dot marks the critical exceptional points where all the lines intersect.
@@
with a similar solution for $\bar{\varphi}_{1}$. This sector, which is plotted for $u\bar{\varphi}_{1}^{2}/|r|\in[0,1/\sqrt{6}]$ in Figure 5, represents one eighth of the total solution set, with the rest being obtained by applying the operations of $\mathrm{D}_{4}$. The critical surface is highlighted in green in Figure 4.
@@
The intersection of the critical surface (highlighted as green in Figure 4) and the exceptional surface (highlighted as purple in Figure 4) is the critical exceptional line, which is shown in red in Figure 4 [see the corresponding equivalent plot for the continuous symmetry case, in Figure 3(d)]. These are located at the intersections of the circle of radius $1/\sqrt{3}$ and the hyperbolas with semi-major axis $\sqrt{\alpha/(3|r|)}$. For $\alpha>|r|$ there are no such points, for $\alpha=0$ and $\alpha=|r|$ there are four, while for $\alpha/|r|\in(0,1)$ there are eight, corresponding to the elements of $\mathrm{D_{4}}$. These observations can be seen in the two-dimensional slices of the phase diagram shown in Figure 6. Additional variations of such plots are presented in the SM[55].
@@ IV Phase separation and NESS
When a mode is unstable, a perturbation of the homogeneous state will grow exponentially, meaning that the system cannot maintain a homogeneous state, and instead undergoes phase separation. The $\alpha=0$ phase diagram in Figure 3 resembles the one-field ($n=1$) double-well Cahn-Hilliard model [48]. In that case, the hashed inner region is stable, and is separated by the spinodal to the metastable region $u\bar{\varphi}^{2}\in(|r|,|r|/3)$ shaded blue. Here, perturbations will not grow exponentially, but a phase-separated system will have lower free energy and is thus thermodynamically favored. In the case of $n=2$, however, this region is unstable, not metastable. The absence of a metastable homogeneous state for the $\mathrm{SO}(2)$-model is a consequence of the set of minima of the potential being connected due to the spontaneously broken continuous symmetry of the system, in contrast to the $\mathrm{C}_{4}$ model, which is phenomenology closer to that of the $n=1$ Cahn-Hilliard. After adding the non-reciprocal term $\alpha$, the corresponding steady-state may no longer be an equilibrium-like, current-free state, but rather a non-equilibrium steady-state (NESS). Still, we may apply similar analysis for insight into these states. We now discuss the phenomenology of phase separation, in and out of equilibrium, for the two cases.
@@ IV.1 Discrete Symmetry
where $\ell=\sqrt{2/|r|}$ is the domain wall thickness, which results from a trade-off between the surface tension term $\nabla^{4}\varphi$ and the potential term $r\nabla^{2}\varphi$. This domain wall profile is, in fact, valid for any average composition $\bar{\varphi}$. As discussed in the previous chapter, there may be more than two phases, but the walls between them will still have a hyperbolic tangent profile with the same length scale $\ell$.
For $\alpha\neq 0$, the considerations become more complicated. As demonstrated recently, the central part of the phase diagram gives rise to traveling wave states [43]. However, at least where $\alpha$ is not too large, the logic of a trade-off between potential and gradient energy—which gives rise to the finite domain wall—still applies. As illustrated in Figure 7, numerical simulations show that states with small $\alpha$ and $\bar{\varphi}$ still form separate phases, whose traveling domain walls are well approximated by a hyperbolic tangent with a thickness $\ell$. In fact, these traveling states may be fully characterized in the thin-wall, low-activity regime, in which their velocity is proportional to the activity and inversely proportional to their size. See video_1.mp4 in[55] for further illustration.

### v2

@@ Abstract
Interactions between active particles may be non-reciprocal, breaking action-reaction symmetry and leading to novel physics not seen in equilibrium systems. In this work, we study the state diagram for the non-reciprocal Cahn-Hilliard (NRCH) model with different types of symmetry—discrete and continuous—and examine the effects of non-reciprocity on phase separation. The models are analyzed by drawing the full linear stability diagrams, indicating the onset of phase separation, allowing us to demonstrate how non-reciprocity gives rise to out-of-equilibrium steady states with non-zero currents. These observations remain valid in a large part of parameter space, and with the addition of weak noise. By describing the effects of symmetry on conserved, non-reciprocal models, this work adds to our understanding of the rich behavior of such models.
@@ II The Model
where summation is implied. The total free energy reads $F=\int\mathrm{d}\bm{x}\,\mathcal{F}$. The two models are defined by their potential $\mathcal{V}(\varphi)$. The potential of the model with $\mathrm{SO}(2)$ symmetry is
@@
If $r\geq 0$, the potential has one global minimum at $\varphi_{a}=0$, whereas for $r<0$ the potential takes the characteristic Mexican hat shape illustrated in Figure 1. The potential is a function of $\varphi_{a}\varphi_{a}$ only, which is a construction that is invariant under rotations and reflections (e.g. $\varphi_{1}\rightarrow-\varphi_{1}$) in the field space. These transformations form the Lie group $\mathrm{O}(2)$. The dynamics arising from this potential corresponds to Model B in the classification of Halperin and Hohenberg [53]. The symmetry is reduced to $\mathrm{SO}(2)$ with the introduction of a non-reciprocal term $\alpha\epsilon_{ab}\varphi_{b}$ in the equations of motion (see below). This reduction cannot happen at the potential level, without also breaking rotational invariance, which is why the potential has a larger symmetry group than the model.
@@
which is not invariant under continuous rotations, as illustrated in Figure 1. However, $\mathcal{V}_{\mathrm{ds}}$ is invariant under operations in $\mathrm{D}_{4}$, the dihedral group of order 8 corresponding to the symmetries of a square. These operations are rotations with $n\pi/2$ radians in addition to reflections, e.g. $\varphi_{1}\rightarrow-\varphi_{1}$. This is a discrete group, similar to the symmetry group of the single-field Cahn-Hilliard $\mathrm{O}(1)\cong\mathrm{D}_{2}$, and in the same way gives rise to a phase separation with a well-defined domain wall. The introduction of a non-reciprocal term restricts the symmetry to $\mathrm{C}_{4}$, namely, the cyclic group of order 4.
Figure 2: The state diagram of the model with continuous symmetry. The exceptional plane is highlighted in purple, and the planes between the stable, type-A unstable, and type-B unstable are highlighted in black, blue, and green, respectively. The red line is the intersection of all these surfaces.
@@ III Linear Stability
We now construct linear stability diagrams for the two cases that represent different symmetries by using linear stability analysis. We study the response of the system to small perturbations around a homogeneous state, represented by $\bar{\varphi}_{a}$. Inserting $\varphi_{a}(\bm{x},t)=\bar{\varphi}_{a}+\delta\varphi_{a}(\bm{x},t)$ into the equations of motion, we obtain the leading order contribution in Fourier space as follows
@@
A non-zero value of $\bar{\varphi}_{a}$ seemingly breaks the assumed symmetry of the model. However, these are still present in the form of rotation of both $\bar{\varphi}$ and $\delta\varphi$. In the case of the continuous symmetry the Ward-Takashi identities related to rotation still apply [55], just as they do when this symmetry is broken spontaneously. The conservation law ensures there are two zero-modes, regardless of whether the continuous symmetry is broken or not, making Goldstone’s theorem non-applicable. In this case, the continuous nature of the symmetry ensures that a domain wall has a divergent length-scale, rather than the correlation length.
@@ III.1 Continuous Symmetry
The base of this diagram is the $\alpha=0$ surface. We choose $\alpha>0$, without loss of generality. At $\alpha=u\bar{\varphi}^{2}$, we have $\xi^{-2}_{+}=\xi^{-2}_{-}$. In reciprocal systems, this indicates a degeneracy where two independent eigenmodes have the same eigenvalue. When we deal with non-reciprocal systems, and thus non-Hermitian matrices, we are no longer guaranteed that the eigenvectors of $M_{ab}$ span the full vector space of perturbations. At this exceptional point (EP) [36, 57], the eigenvalues coincide because the two eigenmodes align, rendering the matrix $M_{ab}$ defective. As there are three free parameters and the exceptional points are given by one equation, namely, $\alpha=u\bar{\varphi}^{2}$, the full set of exceptional points forms a 2D surface, which we denote as the exceptional plane. This is illustrated in purple in the Figure 2. Above the exceptional plane, the two eigenvalues become complex, and the perturbations will no longer decay or diverge monotonically: they exhibit oscillatory behavior. Here, we have $\mathrm{Re}\left(\xi_{\pm}^{-2}\right)=r+2u\bar{\varphi}^{2}$, and, therefore, both modes transition from stable to unstable at the same surface, defined via $r=-2u\bar{\varphi}^{2}$. The intersection between the critical and the exceptional planes can be called critical exceptional points (CEPs) [45]. In the dynamical systems literature, these are know as Takens-Bogdanov bifurcations [58]. These CEPs are parameterized as follows
@@
Figure 3: The four non-trivial slices of the full three-dimensional state diagram shown in Figure 2. Panel (a) corresponds to the equilibrium case $\alpha=0$, while panels (b)-(d) illustrate three orthogonal slices for finite $\alpha$, at constant $\alpha$, $\bar{\varphi}$, and $r$, respectively. Light blue and hashed areas mark the type-A and type-B unstable regions. Black, blue dash-dotted, and green lines separate stable and type-A, type-A and type-B, and stable and type-B, respectively. The purple dashed line marks the exceptional points, and the red dot marks the critical exceptional points where all the lines intersect.
To better present the linear stability diagram, we now examine various two-dimensional slices that can be obtained by fixing one of the parameters. Setting $\alpha=0$ gives the diagram of Model B for $n=2$ fields, with $\xi_{\pm}^{-2}=r+(2\pm 1)u\bar{\varphi}^{2}$. This is illustrated in Figure 3(a), and is the base of the diagram in Figure 2. At $r=\bar{\varphi}=0$, the correlation lengths diverge. This is a critical point corresponding to the onset of phase separation. Starting in the white region of the diagram, where $u\bar{\varphi}^{2}>|r|$, we have $\xi_{\pm}^{-2}>0$, and thus both modes are stable. Upon further lowering $\bar{\varphi}$, we reach a line in the diagram at $u\bar{\varphi}^{2}=|r|$—shown as a black line in Figure 2—where $\xi^{-2}_{-}=0$ and the corresponding eigenmode transitions from stable to unstable. Therefore, perturbations in this direction in $\varphi$-space ($\delta\varphi_{2}$ in this case) will grow exponentially.
Looking back at the Mexican hat potential illustrated in Figure 1, this corresponds to a point inside the ring of minima, and perturbations that would correspond to the Goldstone mode in a spontaneous symmetry breaking scenario (i.e. the orthogonal to the radial direction). The mode corresponding to perturbations in the radial direction ($\delta\varphi_{1}$ in this case) remains stable. This holds in the light blue area labeled “type-A unstable”, where there is one unstable mode. At $u\bar{\varphi}^{2}=|r|/3$, illustrated by a blue dash-dotted line, the second mode also transitions from stable to unstable. The area where both modes are unstable is highlighted as dark blue and hashed, with the label “type-B unstable”.
@@
Figure 4: The state diagram of the model with discrete symmetry, scaled by $|r|$. Light blue and hashed areas mark the type-A and type-B unstable regions. Black, blue, and green planes separate stable and type-A, type-A and type-B, and stable and type-B regions, respectively. The exceptional plane is highlighted in purple, and the red line marks the critical exceptional points where all the planes intersect.
@@ III.2 Discrete Symmetry
Since the eigenvalues depend on both $\bar{\varphi}_{1}^{2}$ and $\bar{\varphi}_{2}^{2}$, the full linear stability diagram requires another dimension to visualize. For this model, we focus on a 3D slice of the full 4D diagram as shown in Figure 4. This cross section is related to any other 3D slice via a choice of scaling, which in this case is chosen to be $|r|$. The inverse-squared correlation lengths defined via $\xi_{\pm}^{-2}=\lambda_{\pm}(q\rightarrow 0)$ as above are then given as follows
@@
Figure 6: Constant $\alpha$ slices of the linear stability diagram illustrated in Figure 4. The gray dashed lines represent the line connecting the minima of the potential illustrated in Figure 1(a), and form a square centered at the origin and with side lengths $\sqrt{u/|r|}\bar{\varphi}=\sqrt{2}$. Light blue and hashed areas mark the type-A and type-B unstable regions. Black, blue dash-dotted, and green lines separate stable and type-A, type-A and type-B, and stable and type-B, respectively. The purple line marks the exceptional points, and the red dot marks the critical exceptional points where all the lines intersect.
@@
with a similar solution for $\bar{\varphi}_{1}$. This sector, which is plotted for $u\bar{\varphi}_{1}^{2}/|r|\in[0,1/\sqrt{6}]$ in Figure 5, represents one eighth of the total solution set, with the rest being obtained by applying the operations of $\mathrm{D}_{4}$. The critical surface is highlighted in green in Figure 4. For a comparison with the continuous case, Figure 4 corresponds to Figure 3(d).
@@
The intersection of the critical surface (highlighted as green in Figure 4) and the exceptional surface (highlighted as purple in Figure 4) is the critical exceptional line, which is shown in red in Figure 4 [see the corresponding equivalent plot for the continuous symmetry case, in Figure 3(d)]. These are located at the intersections of the circle of radius $1/\sqrt{3}$ and the hyperbolas with semi-major axis $\sqrt{\alpha/(3|r|)}$. For $\alpha>|r|$ there are no such points, for $\alpha=0$ and $\alpha=|r|$ there are four, while for $\alpha/|r|\in(0,1)$ there are eight, corresponding to the elements of $\mathrm{D_{4}}$. These observations can be seen in the two-dimensional slices of the linear stability diagram shown in Figure 6. Additional variations of such plots are presented in the SM[56].
@@ IV Phase separation and NESS
When a mode is unstable, a perturbation of the homogeneous state will grow exponentially, meaning that the system cannot maintain a homogeneous state, and instead undergoes phase separation. The $\alpha=0$ linear stability diagram in Figure 3 resembles the one-field ($n=1$) double-well Cahn-Hilliard model [48]. In that case, the hashed inner region is unstable, and is separated by the spinodal to the metastable region $u\bar{\varphi}^{2}\in(|r|,|r|/3)$ shaded blue. Here, perturbations will not grow exponentially, but a phase-separated system will have lower free energy and is thus thermodynamically favored. In the case of $n=2$, however, this region is unstable, not metastable. The absence of a metastable homogeneous state for the $\mathrm{SO}(2)$-model is a consequence of the set of minima of the potential being connected due to the spontaneously broken continuous symmetry of the system, in contrast to the $\mathrm{C}_{4}$ model, which is phenomenology closer to that of the $n=1$ Cahn-Hilliard. After adding the non-reciprocal term $\alpha$, the corresponding steady-state may no longer be an equilibrium-like, current-free state, but rather a non-equilibrium steady-state (NESS). Still, we may apply similar analysis for insight into these states. We now discuss the phenomenology of phase separation, in and out of equilibrium, for the two cases.
@@ IV.1 Discrete Symmetry
where $\ell=\sqrt{2/|r|}$ is the domain wall thickness, which results from a trade-off between the surface tension term $\nabla^{4}\varphi$ and the potential term $r\nabla^{2}\varphi$. This domain wall profile is, in fact, valid for any average composition $\bar{\varphi}$. As discussed in the previous chapter, there may be more than two phases, but the walls between them will still have a hyperbolic tangent profile with the same length scale $\ell$. See Ref. [59] for a more complete discussion.
For $\alpha\neq 0$, the considerations become more complicated. As demonstrated recently, the central part of the phase diagram gives rise to traveling wave states [43]. However, at least where $\alpha$ is not too large, the logic of a trade-off between potential and gradient energy—which gives rise to the finite domain wall—still applies. As illustrated in Figure 7, numerical simulations show that states with small $\alpha$ and $\bar{\varphi}$ still form separate phases, whose traveling domain walls are well approximated by a hyperbolic tangent with a thickness $\ell$. See video_1.mp4 in[56] for further illustration.
@@
For some restricted versions of the NRCH, one may construct a “spurious gradient flow” [60, 61]. These methods, however, are not applicable in our case, due to the symmetries chosen. They balance “spurious” pressure and chemical potential across a stationary domain wall; however, this is not available for traveling waves or domain walls. A model similar to those considered in  [60, 61] is obtained by adding a term $-\rho\varphi_{1}\varphi_{2}$ to the $\mathrm{C}_{4}$ model, breaking its symmetry down to $\mathbb{Z}_{2}$. As seen in recent investigations [62], in models with only $\mathbb{Z}_{2}$-symmetry the onset of traveling-wave states happens for finite non-reciprocity [63, 62]. This thus illustrates how symmetries allow for achieving much “cheaper” non-equilibrium steady states; however, an arbitrarily small $\alpha$ still results in traveling states. Similar behavior was observed in Ref. [59].